# arXiv:2603.00283v1[eess.SY] 27 Feb 2026

### v1

@@ 7.1.1 System Dynamics
$A\!=\!\!\begin{bmatrix}0&0.0067&1.34\\ 0&-0.869&43.2\\ 0&0.993&-1.34\end{bmatrix}\!,B\!\mid\!B_{u}\!=\!\left[\begin{array}[]{cc|c}0.169&0.252&0.1061\\ -17.3&-1.58&0\\ -0.169&-0.252&0.1061\end{array}\right]$
@@ 7.1.2 UC-MPC Design
Based on the uncertainty formulation in Section 7.1.1, for any set ${\mathcal{Z}}\in\mathbb{R}^{3}$ (domain for system state $x(t)$), the values $L_{f_{1},{\mathcal{Z}}}=0.36\max_{\alpha\in{\mathcal{Z}}_{3}}\left\lvert\alpha\right\rvert$, $L_{f_{2},{\mathcal{Z}}}=0.36$, $b_{f_{1},{\mathcal{Z}}}=1.44+0.18\max_{\alpha\in{\mathcal{Z}}_{3}}\alpha^{2}$, and $b_{f_{2},{\mathcal{Z}}}=0.18+0.36\max_{\alpha\in{\mathcal{Z}}_{3}}\left\lvert\alpha\right\rvert$ satisfy Assumption 1. Additionally, $b_{w}=3$ satisfies Assumption 1. For the design of the $\mathcal{L}_{1}$}AC controller in 26, 25 and 24, we chose $A_{e}=-10I_{3}$ and parameterized the filter as ${\mathcal{C}}(s)=\frac{k_{f}}{s+k_{f}}I_{2}$, setting the bandwidth for both input channels as $k_{f}=200$. We also set $\gamma_{1}=0.01$ and the estimation sample time $T$ to $10^{-7}$ seconds, which satisfies 29. To achieve individual performance bounds, $\mathrm{T}_{x}^{i}[k]=0.01$ was assigned for all $i,k\in\mathbb{Z}_{1}^{3}$ with $k\neq i$, ensuring compliance with 34. The reference command $r(t)$ was specified as $[9,6.5]$ deg for $t\in(0,7.5)$ seconds and $[0,0]$ deg for $t\in(7.5,15)$ seconds. For the MPC design for the nominal system, the time horizon was set to 0.2 seconds, and the cost function in 49 at time $\tau$ was defined as: $\mathbf{J}(x_{\textup{n}}(\tau),u(\cdot))=\int_{0}^{0.2}100\|Cx(t)-r(\tau+t)\|+\|u(t)\|+100\|\dot{u}(t)\|\,dt$. The term $\|\dot{u}(t)\|$ was included to penalize control inputs with large time derivatives, ensuring a smoother control input trajectory. For simplicity, we did not include terminal constraint in the MPC formulation.
@@ 7.1.3 Performance bounds with different $K_{x}$
For the longitudinal dynamics of the aircraft, the angle of attack is critical to safety, and thus we aim to design a feedback gain $K_{x}$ such that the effect of the unmatched uncertainty $w(t)$ on $\alpha(t)$ is as small as possible. By solving (20) with $C_{z}=[0,0,1;0,0,0;0,0,0]$ and $D_{z}=[0,0;0.1,0;0,0.1]$, the feedback gain of the feedback controller was obtained as $K_{x}=[4.1075,1.6612,16.2228;-3.0950,0.0595,-1.6320]$ with $A_{m}=A+BK_{x}$ being Hurwitz. Here, we specifically designed $D_{z}$ to include small weights on the control input to ensure the $K_{x}$ is not excessively large and thus it results in a reasonable control input. The calculated individual bounds following Algorithm 1 are listed as follows: $\tilde{\rho}=[0.40,1.28,0.11]$, $\tilde{\rho}_{u}=[7.40,2.26]$, and $\rho_{u_{a}}=[4.34;1.65]$. Thus, the constrained domains for the nominal system were designed as $x(t)\in\mathcal{X}_{\textup{n}}\triangleq[-9.60,9.60]\times[-98.72,98.72]\times[-3.89,3.89]$ and $u(t)\in{\mathcal{U}}_{\textup{n}}\triangleq[-15.07,15.07]\times[-18.86,18.86]$. To illustrate the effect of the design of $K_{x}$ on the constraint tightening parameter $\beta$, we selected other $K_{x}$ matrices that make $A_{m}$ Hurwitz and calculated the corresponding performance bounds. The results are presented in Table 1. It shows that when $K_{x}$ leads to a smaller PPG bound $\beta$, $\tilde{\rho}^{3}$ will be smaller, indicating a smaller effect of $w(t)$ on $\alpha(t)$. However, we achieve the decrease in the state at the cost of a larger control bound, $\tilde{\rho}^{1}_{u}$.
@@ 7.1.4 Simulation Results
The simulation results are shown in Figs. 2, 3, 8, 5, 4, 6 and 7. Figs. 2 and 3 show that UC-MPC resulted in an improved tracking performance compared with both MPC and TMPC, achieving a small error between the actual output $y(t)$ and the reference trajectory $r(t)$. Based on Fig. 4, the matched uncertainty estimation from UC-MPC was accurate, and thus the proposed UC-MPC is able to approximately cancel the matched uncertainty $f(t,x(t))$ with the adaptive control input $u_{a}(t)$. According to the zoomed-in view in Fig. 3, the output trajectories from UC-MPC did not precisely follow the reference trajectories in steady-state. The existence of the unmatched uncertainty $w(t)$ results in the oscillations of the closed-loop trajectory in steady-state, at the same frequency. In contrast, the output trajectories under MPC and TMPC were influenced by both the matched uncertainty and unmatched uncertainty, resulting in a larger tracking error in steady-state.
In terms of enforcing constraints, Fig. 5 demonstrates that both TMPC and UC-MPC managed to satisfy all constraints due to the implementation of constraint tightening. On the other hand, the MPC output breached the state constraint boundary as it did not account for uncertainties during optimization without the use of constraint tightening. It is also worth mentioning that TMPC enforced the constraints at the cost of larger steady-state tracking error, as shown in Fig. 3. Fig. 6 displays the trajectories of the actual state $x(t)$ and the nominal state $x_{\textup{n}}(t)$, as well as the error between them, and the derived individual bounds from $\mathcal{L}_{1}$}AC. Under UC-MPC, the actual states remained consistently close to the nominal states, with the error staying below the computed individual bound. Similarly, as shown in Fig. 7, the adaptive control inputs $u_{\textup{a}}(t)$ remained within the theoretical bounds calculated according to 45b.
We report numerical timing results from the simulation study to compare the computational complexity of the proposed approach with tube-MPC and vanilla MPC. For both tube-MPC and vanilla MPC, the optimization problem must be solved at every control step during real-time execution, resulting in an average computation time of $0.03$ seconds per step. In contrast, for UC-MPC, the nominal MPC problem is solved offline. During online execution, the control input is obtained by combining the pre-computed nominal MPC law with the adaptive compensation term $u_{a}$. Therefore, the only online computation arises from evaluating $u_{a}$, which requires propagating the state predictor and updating the uncertainty estimate through the adaptive law. This process takes an average of $3\times 10^{-5}$ seconds per step. These results demonstrate that the online computational burden of UC-MPC is significantly lower than that of tube-MPC, and that the computational overhead introduced by the adaptive component is negligible.
@@ 7.1.5 Scaling of $w(t)$
$A\!=\!\!\begin{bmatrix}0&0.0067&1.34\\ 0&-0.869&43.2\\ 0&0.993&-1.34\end{bmatrix}\!,B\!\mid\!B_{u}\!=\!\!\left[\begin{array}[]{c|cc}0.169&1.5\text{e-}1&7\text{e-}4\\ -17.3&1.5\text{e-}3&-7\text{e-}4\\ -0.169&-1.5\text{e-}3&7.5\text{e-}2\end{array}\right]$
where the control input $u(t)$ only includes $\delta_{e}(t)$ and the unmatched uncertainty includes two channels. Now, we consider the following artificial uncertainties:
@@ 7.2 Soft landing of a spacecraft on an asteroid
Landing a spacecraft on an asteroid presents distinct challenges. First, the asteroid’s shape, density, and resulting gravity field are inherently difficult to characterize accurately. Consequently, the nominal system used for controller design often differs significantly from the actual, unknown system. This discrepancy means that controllers designed to meet specific performance and safety requirements for the nominal system may fail to ensure these properties for the real system. Second, the vast distance between the spacecraft and Earth introduces communication delays, making real-time ground-based control infeasible. As a result, the spacecraft must be made autonomous and made rely entirely on onboard control frameworks. Given these challenges, the UC-MPC framework emerges as a promising solution, offering robust performance and safety guarantees for achieving a soft landing on an asteroid despite significant uncertainties.
@@ 7.2.1 System Dynamics
with all $S_{l,m}$ in (57) are zero and all other $C_{l,m}$ are zero. Using the coefficients calculated in (58), the original nonlinear dynamics (56) can be linearized about a target position equilibrium, e.g., $X_{0}=[x_{0}\;y_{0}\;z_{0}\;0\;0\;0]^{\top}$, as:
@@ 7.2.3 Simulation Results
To validate the robustness of the proposed UC-MPC framework, we conducted additional tests in Phase I using varying initial positions while keeping the target position constant. The results, shown in Fig. 14, demonstrate that the proposed control framework performs reliably with small state error (less than 0.15 km) between the real system and the nominal system across different initial conditions, highlighting its robustness to variations in starting positions.
@@ A.2 Proof of Lemma 4
$=G_{xm}(s)\mathfrak{L}\left[\eta_{\textup{r}}(t)\right]+\mathcal{H}_{xu}(s)w(s).$ (89)

### v3

@@ 7.1.1 System Dynamics
$A\!=\!\!\begin{bmatrix}0&0.0067&1.34\\ 0&-0.869&43.2\\ 0&0.993&-1.34\end{bmatrix}\!,B\!\mid\!B_{u}\!=\!\left[\begin{array}[]{cc|c}0.169&0.252&0.1061\\ -17.3&-1.58&0\\ -0.169&-0.252&0.1061\end{array}\right]\!,$
@@ 7.1.2 UC-MPC Design
Based on the uncertainty formulation in Section 7.1.1, for any set ${\mathcal{Z}}\in\mathbb{R}^{3}$ (domain for system state $x(t)$), the values $L_{f_{1},{\mathcal{Z}}}=0.36\max_{\alpha\in{\mathcal{Z}}_{3}}\left\lvert\alpha\right\rvert$, $L_{f_{2},{\mathcal{Z}}}=0.36$, $b_{f_{1},{\mathcal{Z}}}=1.44+0.18\max_{\alpha\in{\mathcal{Z}}_{3}}\alpha^{2}$ and $b_{f_{2},{\mathcal{Z}}}=0.18+0.36\max_{\alpha\in{\mathcal{Z}}_{3}}\left\lvert\alpha\right\rvert$ satisfy Assumption 1. Additionally, $b_{w}=3$ satisfies Assumption 1. For the design of the $\mathcal{L}_{1}$}AC controller in 26, 25 and 24, we chose $A_{e}=-10I_{3}$ and parameterized the filter as ${\mathcal{C}}(s)=\frac{k_{f}}{s+k_{f}}I_{2}$, setting the bandwidth for both input channels as $k_{f}=200$. We also set $\gamma_{1}=0.01$ and the estimation sample time $T$ to $10^{-7}$ seconds, which satisfies 29. To achieve individual performance bounds, $\mathrm{T}_{x}^{i}[k]=0.01$ was assigned for all $i,k\in\mathbb{Z}_{1}^{3}$ with $k\neq i$, ensuring compliance with 34. The reference command $r(t)$ was specified as $[9,6.5]$ deg for $t\in(0,7.5)$ seconds and $[0,0]$ deg for $t\in(7.5,15)$ seconds. For the MPC design for the nominal system, the time horizon was set to 0.2 seconds, and the cost function in 49 at time $\tau$ was defined as: $\mathbf{J}(x_{\textup{n}}(\tau),u(\cdot))=\int_{0}^{0.2}100\|Cx(t)-r(\tau+t)\|+\|u(t)\|+100\|\dot{u}(t)\|\,dt$. The term $\|\dot{u}(t)\|$ was included to penalize control inputs with large time derivatives, ensuring a smoother control input trajectory. For simplicity, we did not include terminal constraint in the MPC formulation.
@@ 7.1.3 Performance bounds with different $K_{x}$
For the longitudinal dynamics of the aircraft, the angle of attack is critical to safety, and thus we aim to design a feedback gain $K_{x}$ such that the effect of the unmatched uncertainty $w(t)$ on $\alpha(t)$ is as small as possible. By solving (20) with $C_{z}=[0,0,1;0,0,0;0,0,0]$ and $D_{z}=[0,0;0.1,0;0,0.1]$, the feedback gain of the feedback controller was obtained as $K_{x}=[4.1075,1.6612,16.2228;-3.0950,0.0595,-1.6320]$ with $A_{m}=A+BK_{x}$ being Hurwitz. Here, we specifically designed $D_{z}$ to include small weights on the control input to ensure that $K_{x}$ is not excessively large and thus it results in a reasonable control input. The calculated individual bounds following Algorithm 1 are listed as follows: $\tilde{\rho}=[0.40,1.28,0.11]$, $\tilde{\rho}_{u}=[7.40,2.26]$, and $\rho_{u_{a}}=[4.34;1.65]$. Thus, the constrained domains for the nominal system were designed as $x(t)\in\mathcal{X}_{\textup{n}}\triangleq[-9.60,9.60]\times[-98.72,98.72]\times[-3.89,3.89]$ and $u(t)\in{\mathcal{U}}_{\textup{n}}\triangleq[-15.07,15.07]\times[-18.86,18.86]$. To illustrate the effect of the design of $K_{x}$ on the constraint tightening parameter $\beta$, we selected other $K_{x}$ matrices that make $A_{m}$ Hurwitz and calculated the corresponding performance bounds. The results are presented in Table 1. It shows that when $K_{x}$ leads to a smaller PPG bound $\beta$, $\tilde{\rho}^{3}$ will be smaller, indicating a smaller effect of $w(t)$ on $\alpha(t)$. However, we achieve the decrease in the state at the cost of a larger control bound, $\tilde{\rho}^{1}_{u}$.
@@ 7.1.4 Simulation Results
The simulation results are shown in Figs. 2, 3, 8, 5, 4, 6 and 7. First, Figs. 2 and 3 show that UC-MPC resulted in an improved tracking performance compared with both MPC and TMPC, achieving a small error between the actual output $y(t)$ and the reference trajectory $r(t)$. Based on Fig. 4, the matched uncertainty estimation from UC-MPC was accurate, and thus the proposed UC-MPC is able to approximately cancel the matched uncertainty $f(t,x(t))$ with the adaptive control input $u_{a}(t)$. According to the zoomed-in view in Fig. 3, the output trajectories from UC-MPC did not precisely follow the reference trajectories in steady-state. The existence of the unmatched uncertainty $w(t)$ results in the oscillations of the closed-loop trajectory in steady-state, at the same frequency. In contrast, the output trajectories under MPC and TMPC were influenced by both the matched uncertainty and unmatched uncertainty, resulting in a larger tracking error in steady-state.
In terms of enforcing constraints, Fig. 5 demonstrates that both TMPC and UC-MPC managed to satisfy all constraints due to the implementation of constraint tightening. On the other hand, the MPC output breached the state constraint boundary as it did not account for uncertainties during optimization without the use of constraint tightening. It is also worth mentioning that TMPC enforced the constraints at the cost of larger steady-state tracking error, as shown in Fig. 3. Figure 6 displays the trajectories of the actual state $x(t)$ and the nominal state $x_{\textup{n}}(t)$, as well as the error between them, and the derived individual bounds from $\mathcal{L}_{1}$}AC. Under UC-MPC, the actual states remained consistently close to the nominal states, with the error staying below the computed individual bound. Similarly, as shown in Fig. 7, the adaptive control inputs $u_{\textup{a}}(t)$ remained within the theoretical bounds calculated according to 45b.
For tube MPC, solving the optimization problem with the additional decision variable associated with the initial condition results in an average computation time of $0.05$ seconds per control step. In contrast, the optimization problem in UC-MPC has essentially the same structure as a standard MPC formulated purely on the nominal system. As a result, the optimization is simpler and requires an average computation time of $0.03$ seconds per step. The only additional online computation arises from evaluating the adaptive control input $u_{a}$, which involves propagating the state predictor and updating the uncertainty estimate through the adaptive law. This procedure requires an average of only $3\times 10^{-5}$ seconds per step. These results demonstrate that UC-MPC significantly reduces the online computational burden compared with tube MPC, while the computational overhead introduced by the adaptive component is negligible.
@@ 7.1.5 Scaling of $w(t)$
$A\!=\!\!\begin{bmatrix}0&0.0067&1.34\\ 0&-0.869&43.2\\ 0&0.993&-1.34\end{bmatrix}\!,B\!\mid\!B_{u}\!=\!\!\left[\begin{array}[]{c|cc}0.169&1.5\text{e-}1&7\text{e-}4\\ -17.3&1.5\text{e-}3&-7\text{e-}4\\ -0.169&-1.5\text{e-}3&7.5\text{e-}2\end{array}\right]\!,$
where the control input $u(t)$ only includes $\delta_{e}(t)$, and the unmatched uncertainty includes two channels. Now, we consider the following artificial uncertainties:
@@ 7.2 Soft landing of a spacecraft on an asteroid
Landing a spacecraft on an asteroid presents distinct challenges. First, the asteroid’s shape, density, and resulting gravity field are inherently difficult to characterize accurately. Consequently, the nominal system used for controller design often differs significantly from the actual, unknown system. This discrepancy means that controllers designed to meet specific performance and safety requirements for the nominal system may fail to ensure these properties for the real system. Second, the vast distance between the spacecraft and Earth introduces communication delays, making real-time ground-based control infeasible. As a result, the spacecraft must be made autonomous and made to rely entirely on onboard control frameworks. Given these challenges, the UC-MPC framework emerges as a promising solution, offering robust performance and safety guarantees for achieving a soft landing on an asteroid despite significant uncertainties.
@@ 7.2.1 System Dynamics
with all $S_{l,m}$ in (57) being zero and all other $C_{l,m}$ being zero. Using the coefficients calculated in (58), the original nonlinear dynamics (56) can be linearized about a target position equilibrium, e.g., $X_{0}=[x_{0}\;y_{0}\;z_{0}\;0\;0\;0]^{\top}$, as:
@@ 7.2.3 Simulation Results
To validate the robustness of the proposed UC-MPC framework, we conducted additional tests in Phase I using varying initial positions while keeping the target position constant. The results, shown in Fig. 14, demonstrate that the proposed control framework performs reliably with small state error between the real system and the nominal system across different initial conditions, highlighting its robustness to variations in starting positions.
@@ A.2 Proof of Lemma 4
$=G_{xm}(s)\mathfrak{L}\left[\eta_{\textup{r}}(t)\right]+\mathcal{H}_{xu}(s)w(s)$ (89)